If the sum of two unit vectors is a unit vector then find the magnitude of their difference

Grade:11

1 Answers

Arun
6 years ago
Dear Kunal
 

let two vectors are a,b then

 

resultant of sum = (a2 + b2 + 2abcos@)1/2 = 1               (given)

 

a = b = 1          (unit vector)

 

 1+1+2cos@ = 1

 

      cos@ = -1/2           .................1

 

now resultant of difference = magnitude (a-b)

 

R = (a2 + b2 - 2abcos@)1/2

 

a = b = 1 , cos@ = -1/2 so

 

R = [ 1+1+1]1/2 = root3